OpenStudy (anonymous):

Which of the following is the equation of a line that passes through (-2, -1) and (-4, -3)?

OpenStudy (anonymous):

What two things do you need in order to find the equation of a line?

OpenStudy (anonymous):

x and y ?

OpenStudy (anonymous):

no you need a point and a slope. You are given two points so you can pick either. All you have to do now is find the slope and plug that into this equation \[y-y_a=m(x-x_a)\] where\[m=slope \\ (x_a,y_a)= point \]

OpenStudy (anonymous):

Alright do you know how to find a slope between two points?

OpenStudy (anonymous):

I dont remember ..

OpenStudy (anonymous):

\[m=\frac{(y_a-y_b)}{(x_a-x_b)}\] where \[(x_a,y_a)\]\[(x_b,y_b)\] are the the two points given to you. You think you can find the slope now?

OpenStudy (anonymous):

slope is -1

OpenStudy (anonymous):

@russellp i dont think that's right....can you show your work

OpenStudy (anonymous):

(-2 ,-1 ) (-4 , -3) y2 -y1 / x2 -x1 = slope -4 - -1 / -3 - -1 = 5/4 = 1.25 corrected sorry Does that calculation seem correct?

OpenStudy (anonymous):

you mixed the y and x together as well as subtracting wrong

OpenStudy (anonymous):

-3 - -1/ -4 - -2 = -2 / -2 = 1 slope ? Try this one ???? Thank you

OpenStudy (anonymous):

The answer is 1 ?

OpenStudy (anonymous):

How do i make that into an equation then ?

OpenStudy (anonymous):

the equation for the line is y=mx + B where m is the slope of the line and b is the y intercept therefore calculate from there

OpenStudy (anonymous):

Thank you